Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.50 (given its target capital structure). Vandell has $9.62 million in debt that trades at par and pays an 8% interest rate. Vandell’s free cash flow (FCF0) is $2 million per year and is expected to grow at a constant rate of 5% a year. Both Vandell and Hastings pay a 30% combined federal and state tax rate. The risk-free rate of interest is 5% and the market risk premium is 5%. Hastings Corporation estimates that if it acquires Vandell Corporation, synergies will cause Vandell’s free cash flows to be $2.6 million, $2.7 million, $3.4 million, and $3.60 million at Years 1 through 4, respectively, after which the free cash flows will grow at a constant 5% rate. Hastings plans to assume Vandell’s $9.62 million in debt (which has an 8% interest rate) and raise additional debt financing at the time of the acquisition. Hastings estimates that interest payments will be $1.6 million each year for Years 1, 2, and 3. After Year 3, a target capital structure of 30% debt will be maintained. Interest at Year 4 will be $1.426 million, after which the interest and the tax shield will grow at 5%. Indicate the range of possible prices that Hastings could bid for each share of Vandell common stock in an acquisition. Round your answers to the nearest cent. Do not round intermediate calculations. The bid for each share should range between
Answer #1

a. Compute cost of equity and WACC

Beta =1.4
Risk free rate =5%   
Market risk =6%

Cost of equity= Risk free rate+Market risk*Beta
=5+6*1.4
=5+8.4
=13.40%


Before tax cost of debit=8%
Tax rate=40%
After tax cost of debt=(1-Tax rate)*Before tax cost of debit
=(1-40/100)*8/100
=(1-0.4)*0.08
=0.6*0.08
=0.048


Weight of debt=0.3=30%
Weight of equity =0.7=70%
WACC= Weight of debt*After tax cost of debt+Weight of equity*Cost of equity
=0.3*0.048+0.7*0.0134
=0.02378
=2.378%

b. Compute value of operation
------------------------------------------
FCF0 =2000000
Constant growth rate=0.05
Value of operation=(1+Constant growth rate)*FCF/(Constant growth rate-WACC)

=(1+0.05)*2000000/(0.05-0.02378)
=77,777,777.777

  
C.Compute intrisic value per share
_____________________________
Value of debt =10820000
Value of equity=Value of operation-Value of debt
=77,777,777.77-10820000
=66,957,777.77
Number of shares= 1000000
Intrinsic per share=Value of equity/Number of shares
=66.95